OpenStudy (anonymous):

Explain each step. You must use the terms reciprocal and coefficient at least once. Label your steps with Step 1, Step 2, Step 3, and Step 4. Equation Justification two-thirds times x equals four-fifths the reciprocal of two-thirds is three-halves 1. three-halves times two-thirds times x equals three-halves times four-fifths 2. x equals three-halves times four-fifths equals twelve-tenths 3. x equals twelve divided by two over ten divided by two, which equals six-fifths which equals one remainder one which equals one and one-fifth 4. will give medal and fan for helping me

OpenStudy (anonymous):

please help me.

OpenStudy (anonymous):

really need either a screen shot or the use of the equation tool. this is why math uses symbols and not words (unless you are writing a check!)

OpenStudy (anonymous):

2/3 * x = 4/5 Multiply both sides by reciprocal of 2/3, which is 3/2 3/2 * 2/3 * x = 3/2 * 4/5 Simplify: x = 3/1 * 2/5 Multiply: x = 6/5 = 1 ¹/₅ −4d/3 = 3/25 Multiply both sides by 3 −4d = 9/25 Divide both sides by coefficient of d (−4) d = −9/100 −8/9 * c = −24 c = 24 * 9/8 c = 3 * 9/1 c = 27 3/4 * a = 12 a = 12 * 4/3 a = 16